OpenStudy (anonymous):

Can someone help me convert this (f(x)=9.75e^-0.43x) into a logarithmic function? Please.

OpenStudy (anonymous):

Suppp.

OpenStudy (anonymous):

.... >.>

OpenStudy (anonymous):

Yeah bro what's up. lol

OpenStudy (anonymous):

I'm not a bro but the question is what's up there. See what I did there? lol.

OpenStudy (amistre64):

log each side ...

OpenStudy (anonymous):

This what I have so far: y=9.75e^(-0.43x) 4/9.75= e^(-0.43x) In the example I have it says something about IN and I have no clue what that is.

OpenStudy (amistre64):

LN, is an undergradute notation for a natural logarithm, in upper classes its just LOG, which i admit can be confusing

OpenStudy (amistre64):

y=9.75e^(-0.43x) ln(y) = ln(9.75e^(-0.43x)) ln(y) = ln(9.75) + ln(e^(-0.43x)) ln(y) = ln(9.75) -0.43x ln(e) ln(y) = ln(9.75) -0.43x ln(y)/ln(9.75) = -0.43x

OpenStudy (anonymous):

Thank you so much.

OpenStudy (amistre64):

yep ... good luck ;)

OpenStudy (anonymous):

@amistre64 check out this law of logarithms: \[\ln(a) - \ln(b) =\ln( \frac{ a}{ b })\]

OpenStudy (amistre64):

yes, and ln(ab) = ln(a) + ln(b) ...

OpenStudy (amistre64):

...and it looks like a divided instead of subtracting; i blame the ascii :)

OpenStudy (amistre64):

ln(y) = ln(9.75) -0.43x ln(y) - ln(9.75) = -0.43x ln(9.75) - ln(y) = 0.43x